OpenStudy (anonymous):

integral of (cosx+sqrt(sins))^2

OpenStudy (anonymous):

my first guess would be to multiply out

OpenStudy (anonymous):

try \[\int \sin(x)+\cos^2(x)+2\sqrt{\sin(x)} \cos(x)dx\]

OpenStudy (anonymous):

first of obvious last is a simple u - sub middle is one of those double angle identities

OpenStudy (anonymous):

by which i mean \[\int \cos^2(x)dx=\frac{1}{2}\int (1+\cos(2x))dx\]